If you dropped a ball while standing on the surface of the sun at what rate would it accelerate toward the ground
1
Expert's answer
2020-10-05T10:56:15-0400

The expression for the gravitational acceleration is:


"g = G\dfrac{M}{R^2}"

where "G = 6.674\times10^{-11} N\cdot m^2/kg^2" is the gravitational constant, "M = 1.989\times 10^{30}kg" is the mass of the Sun and "R = 0.7\times 10^{9} m" is the radius of the Sun. Thus, obtain:


"g = G\dfrac{M}{R^2} = 6.674\times10^{-11} \dfrac{1.989\times 10^{30}}{ 0.7^2\times 10^{18}} \approx 270.9m/s^2"

Answer. 270.9 m/s^2.